The investor owns 40% ownership in the investee. Suppose the investee sells merchandise costing $40,000 to the investor for $60,000, and at year’s end, the investor still retains $15,000 of the goods. The investee reports a net income of $120,000 for the year. The investor uses the equity method.

Prepare journal entries for the investor at the end of the year.
Answer #1

GAAP recognises three method of financial reporting of Equity investment  

1) Fair Value

2) Consolidation of Financial statement

3) Equity method

Equity method employs the accrual basis for recognising the investor's share of Investee income , Investor recognized increase in investment value according to their ownership in investee.

Under Equity method ,investment is initially recognise at cost and get increase or decrease as per the investor share in profit .

Equity accounting is usually applied where an investor entity holds 20–50% of the voting stock of the associate company

Entry for Recognising investor Share in Investee profit

1) Entry will be passed for recognising gross profit Share

Debit Credit
Equity Investment ( $120000×40%) $48000
To Equity Income $48000
(Te recognised increase in Equity investmemt)

Entry to deferred unrealised gain on ending inventory on intercompany sale

2)

Debit Credit
Equity Income 5000×40% 2000
To Equity Investment 2000
( To deffered unrealised gain on investor Ending inventory )

Unrealised gain lying on the Ending inventory of Investor which is included in Profit of Investee

=15000×(60000-40000)/60000

=5000

Portion of unrealised gain on sale included in Investee net income= 5000×40% =2000

Profit margin on sale of Investee to investor is 20000/60000 =33.33333%

Above entry will get reverse on recognition of gain lying on the ending inventory of the investor